# Balancing Accuracy and Efficiency for Large-Scale SLAM: A Minimal Subset Approach for Scalable Loop Closures

### v2

@@ II-E3 Dynamic Keyframe Selection
To minimize the keyframe set size in real-time, we must consider the entire set $\mathbf{K}$; however, this is impractical for maintaining computational feasibility in loop closure detection. The challenge is to determine the contribution of each keyframe, which requires anticipating future keyframes and introduces a lack of causality. Moreover, finding the optimal subset $\mathbf{K}_{\text{S}}^{*}$ is an NP-hard problem, involving an exhaustive search of all keyframe combinations, which becomes computationally infeasible for sets larger than 15–20 keyframes.
The goal of this research is to show that reducing the number of keyframes through an informed sampling strategy can significantly decrease the search space for loop closure detection, improving computational efficiency by reducing memory usage and query time without compromising SLAM accuracy. The proposed approach seeks to balance computational complexity and optimization precision, ensuring scalability and robustness.
III Minimal Subset Approach
@@
To address these challenges, we propose the Minimal Subset Approach (MSA) to approximate a solution to the problems in Eq. (5)-(6). Unlike our previous work [6], which focuses solely on place recognition retrieval performance, in this article we extend the analysis to assess its impact on the subsequent step described in Eq.(7) and its comparison to the original Eq. (2). The proposed MSA identifies and eliminates redundancy within a keyframe set while preserving essential data. Designed for real-time use, it employs a sliding window combinatorial optimization with dual objectives: balancing information preservation and redundancy minimization.
III-A Redundancy in Keyframes
@@
¯=$\bar{\mathcal{L}}\equiv\mathcal{L}=\big{\{}\langle i,j\rangle\mid f(\mathbf{d}_{i},\mathbf{d}_{j})>\tau,\;\forall\langle\mathbf{k}_{i},\mathbf{k}_{j}\rangle\in\bar{\mathbf{K}}\times\bar{\mathbf{K}}\big{\}},$
@@
where $\bar{\mathbf{K}}=\mathbf{K}\,\backslash\{\mathbf{k}\}$. The spatial constraint in Eq. (8) ensures comprehensive map coverage and that consecutive odometry edges can be computed safely. The lower and upper pose distances, $\delta_{l}$ and $\delta_{u}$, usually range from 1 to 5 meters [23]. We quantify redundancy in a keyframe set using a metric that captures the similarity between consecutive keyframes, utilizing any similarity function $f_{\sigma}$, provided by the corresponding descriptor extraction framework:
@@
III-B Information Preservation in Keyframes
@@
Let $F:\mathbb{R}^{3}\rightarrow\mathbb{R}^{M}$, represent the descriptor extraction function, either learning-based [23, 28] or handcrafted [24, 29], which maps each observation from 3-D space to an $M$-dimensional representation. This function does not directly depend on the pose $\mathbf{x}_{t}$, but rather on the input observation $\mathbf{z}_{t}$ (e.g. LiDAR scan), which itself depends on the pose through the sensor model $\mathbf{z}_{t}\propto h(\mathbf{x}_{t},\mathbf{W})$, where $\mathbf{W}$ represents the world. To understand how the descriptors are sensitive to pose changes, we can compute the Jacobian $\mathbf{J}$ of $F$ with respect to poses $\mathbf{x}$ through the chain rule $(\partial\mathbf{F}\,\partial\mathbf{z})/(\partial\mathbf{z}\,\partial\mathbf{x})$. Explicitly deriving the Jacobian can be impractical because of the intricate nature of functions like deep neural networks. Consequently, we employ numerical approximations to estimate the rate of change. Moreover, since poses are encompassed within $SE(3)$, we reduce the dimensional complexity by using the Euclidean norm for distance measurement instead of individual axis derivatives. It is essential to use yaw-invariant descriptors, as suggested in [23, 24], to ensure the Jacobian calculations are meaningful by preventing orientation changes. For a comprehensive discussion of these assumptions and derivations, we refer the reader to our previous work [6], which provides detailed explanations. Considering the descriptors $\mathbf{d}\in\mathbb{R}^{M}$ as random variables and poses $\mathbf{x}\in SE(3)$ as samples, the product $\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}$ estimates the covariance matrix:
$\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}=\bigg{(}\frac{\partial\mathbf{F}}{\partial\mathbf{x}}\bigg{)}^{\mathsf{T}}\bigg{(}\frac{\partial\mathbf{F}}{\partial\mathbf{x}}\bigg{)}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}\raisebox{4.95134pt}{$\scriptscriptstyle-\!1$},$ (10)
where $\mathbf{\Lambda}$ is the diagonal matrix of eigenvalues and $\mathbf{V}$ is the matrix of eigenvectors from the decomposition. The eigenvectors represent the principal directions of variation, and the eigenvalues represent their magnitudes. The descriptors are transformed using the eigenvectors and scaled by the square root of the eigenvalues, $\mathbf{D}^{\prime}=\sqrt{\mathbf{\Lambda}}\cdot\mathbf{V}\cdot\mathbf{D}$, aligning them with the main directions of variability. Therefore, we can define the information preservation term for a keyframe set as:
$\pi_{\tau}(\mathbf{K})=-\frac{1}{N}\sum_{i=0}^{N-1}f_{\delta}(\mathbf{d}_{i}^{\prime},\mathbf{d}_{i+1}^{\prime}),$ (11)
@@
where $-1\leqslant\pi_{\tau}(\mathbf{K})<0$, $f_{\delta}$ is the distance function between descriptors, and $\mathbf{d}^{\prime}\in\mathbf{D}^{\prime}$ are the transformed descriptors. Higher values of $\pi_{\tau}$ indicate that the poses better preserve the variability in the descriptor space. Eigenvectors $\mathbf{v}_{k}$ denote directions of maximal variability, and the $k$-th eigenvalue $\lambda_{k}$ quantifies the variance described by each eigenvector. Larger eigenvalues indicate more significant patterns of variability.
@@
Figure 2: Redundant keyframes effect. Example of removing redundant keyframes (light blue). The updated odometry edge set, $\mathcal{E}_{o}^{*}$, influences the pose graph optimization, while the sparsified loop closure edges, ¯$\bar{\mathcal{L}}=\mathcal{E}_{l}\backslash\langle 5,7\rangle$, can further reduce the computational complexity.
III-C Sliding Window Optimization
To manage the computational complexity of optimizing a large keyframe set and proactively selecting keyframes for future queries, we introduce a sliding window optimization method. This approach continuously optimizes keyframe sets $\mathbf{K}_{t}$ over the mission duration $T$, approximating the optimal keyframe set as:
@@ III-C Sliding Window Optimization
The process begins by initializing a window keyframe set $\mathbf{K}_{t}$ containing $N$ keyframes. The time step $t$ advances once $N$ new keyframes are available, and the optimization converges to the optimal window keyframe set $\mathbf{K}_{t}^{*}$. To find this optimum, we generate all possible subsets of $\mathbf{K}_{t}$, forming the power set $\mathbb{P}(\mathbf{K}_{t})$ with cardinality $|\mathbb{P}(\mathbf{K}_{t})|=2^{N}$. To reduce computational complexity, we apply constraints on the power set, retaining only subsets that satisfy minimum and maximum distances between consecutive poses, as per Eq. (8), reducing the power set by 5–10 times. The constrained power set can be denoted as:
¯$\bar{\mathbb{P}}(\mathbf{K}_{t})=\big{\{}\mathbf{K}_{t}^{\mathbb{S}}\in\mathbb{P}(\mathbf{K}_{t})|\delta_{l}\leqslant\|\mathbf{x}_{i}-\mathbf{x}_{i+1}\|_{2}\leqslant\delta_{u},\forall\mathbf{k}\in\mathbf{K}_{t}^{\mathbb{S}}\big{\}},$ (13)
where $\delta_{l}$ and $\delta_{u}$ are the lower and upper limits for the distance between poses. The window optimization problem can be formulated as:
$\mathbf{K}_{t}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{K}_{t}^{\mathbb{S}}}\>\>\big{(}{\rho_{\tau}\,(\mathbf{K}_{t}^{\mathbb{S}})+\alpha}\big{)}\big{/}\big{(}{\pi_{\tau}(\mathbf{K}_{t}^{\mathbb{S}})}-\beta\big{)}$ (14)
@@
The minimization problem is solved through an exhaustive search, evaluating each subset’s information matrix $\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}$ and quantifying its effectiveness using the information preservation and redundancy terms, selecting the subset with the best combined score.
IV Experimental Setup and Results
TABLE I: Translational (t) and Rotational (R) Absolute Trajectory Error Improvement after PGO and the False Positive Rate (FPR)
@@ IV Experimental Setup and Results
Metrics t. ATE [%] R. ATE [%] FPR [%] t. ATE [%] R. ATE [%] FPR [%] t. ATE [%] R. ATE [%] FPR [%]
ScanContext All Samples 72.7 / 76.1 94.4 / 49.0 06.6 / 06.4 30.2 / 69.4 24.1 / 56.3 08.8 / 20.8 92.0 / 93.7 93.7 / 92.3 29.4 / 06.8
Const. 1.0m 90.5 / 76.2 95.9 / 52.6 04.3 / 04.1 63.2 / 88.6 55.2 / 70.7 11.6 / 24.0 84.3 / 92.5 87.4 / 92.0 26.1 / 05.5
Const. 2.0m 86.2 / 62.3 94.0 / 46.4 06.0 / 05.9 72.1 / 83.3 69.4 / 72.0 12.1 / 13.3 66.7 / 79.2 59.5 / 78.6 25.9 / 03.3
Const. 3.0m 79.8 / 60.1 77.4 / 42.3 10.5 / 09.9 58.7 / 79.5 60.6 / 69.2 19.6 / 11.7 54.4 / 77.1 40.9 / 69.8 24.0 / 03.1
Entropy-based 66.9 / 46.7 93.4 / 48.3 09.1 / 05.5 68.7 / 71.8 69.0 / 66.8 14.2 / 15.6 90.1 / 80.3 89.8 / 82.9 24.5 / 03.2
Spaciousness 77.8 / 55.6 87.6 / 46.7 09.3 / 05.9 67.9 / 70.7 68.5 / 66.1 10.0 / 16.5 79.7 / 79.6 84.2 / 82.1 25.2 / 06.0
MSA (Ours) 90.3 / 76.4 96.1 / 53.0 03.3 / 03.4 71.9 / 91.9 70.1 / 76.6 06.8 / 12.3 89.9 / 87.5 91.4 / 80.3 23.3 / 02.7
OverlapTrans All Samples 75.7 / 78.6 95.9 / 51.1 05.1 / 05.6 29.8 / 67.4 24.9 / 59.5 08.0 / 21.2 96.2 / 98.1 95.3 / 97.5 30.7 / 07.5
Const. 1.0m 94.6 / 78.7 97.3 / 55.3 06.2 / 07.0 59.3 / 87.1 56.4 / 71.0 20.2 / 23.8 85.8 / 97.1 84.1 / 97.2 25.1 / 04.1
Const. 2.0m 84.4 / 44.9 97.2 / 23.7 04.3 / 03.1 70.0 / 80.9 68.4 / 73.7 07.6 / 11.6 06.8 / 75.9 06.8 / 77.2 25.3 / 02.3
Const. 3.0m 90.5 / 22.6 87.9 / 13.8 06.0 / 09.5 63.6 / 87.7 64.6 / 79.1 04.4 / 07.7 00.0 / 83.6 00.0 / 84.4 16.6 / 02.4
Entropy-based 67.9 / 24.1 93.0 / 09.1 06.1 / 05.4 69.2 / 68.7 69.5 / 63.1 16.4 / 14.6 77.6 / 69.4 76.2 / 72.6 21.0 / 03.1
Spaciousness 85.8 / 35.5 97.2 / 39.1 06.8 / 06.9 68.4 / 68.6 68.1 / 63.1 08.5 / 16.1 03.5 / 67.9 03.5 / 71.7 24.1 / 04.9
MSA (Ours) 93.2 / 78.8 97.5 / 55.5 04.0 / 04.9 70.1 / 91.4 69.7 / 77.8 08.2 / 11.4 88.7 / 84.0 88.3 / 85.1 21.5 / 01.6
@@
The underlined results correspond to the best performing constant sampling interval, while the bold results correspond to the best adaptive sampling method.
@@
To support the claims of our second contribution, on how the sampling interval affects performance, we conducted a series of experiments with the following setup. Descriptor extraction was performed using two approaches: the learning-based OverlapTransformer (OT[23], which generates a $1\times 256$ feature vector, and the hand-crafted Scan Context (SC[24], which produces a $20\times 60$ feature representation. These choices demonstrate the descriptor-agnostic nature of our method and its compatibility with both learned and hand-crafted descriptors. We benchmark our approach against constant sampling intervals of 1, 2, and 3 meters [15, 17, 16, 30], as well as adaptive sampling strategies based on LiDAR spaciousness [18] and entropy [20]. All experiments were conducted on a 14th Gen Intel Core i9 system with 128GB of RAM, where the sampling window optimization ($N=10$) is solved in approximately 5–15 milliseconds, making it suitable for real-time sampling.
@@
IV-A Quantitative and Qualitative Analysis
For the batch experiments in Table I, we evaluate our approach on three datasets: KITTI Odometry [31], MulRan [32], and Apollo-SouthBay [33], representing diverse environments, including urban areas, rural landscapes, and complex structures like bridges and tunnels. Keyframes are sampled for all methods, and the pose graph is constructed using GTSAM [7] with odometry provided from KISS-ICP [34]. Loop closure candidates are identified using OT or SC descriptors with a similarity threshold of 0.8. Candidates within a 3-meter radius of the ground truth are classified as true positives and processed with small_gicp [35] to estimate relative transformations. Matches with registration residuals below 0.3 meters are verified and added as edges to the pose graph, which is optimized using the Levenberg-Marquardt algorithm.
Table I summarizes the results for all datasets, descriptors, and methods. The metrics include the Absolute Trajectory Error (ATE) for both translation and rotation, presented as the percent improvement in the KISS-ICP trajectory after pose graph optimization with each method. Additionally, the False Positive Ratio (FPR) for descriptor matching is reported, as sampling impacts place recognition performance [6]. The results demonstrate that the proposed MSA consistently delivers the best performance, while constant intervals yield varying results across different environments, highlighting the absence of a universal fixed sampling interval. In Fig. 3, the data from Table I are presented as box plots, illustrating the deviation of each method’s ATE performance compared to using all samples, alongside variations in memory usage and total pipeline execution time. The proposed approach sustains overall improvements with minimal performance loss in both translation and rotation, while significantly reducing memory usage and execution time. Although other methods also reduce memory and processing overhead, their performance lacks robustness and varies considerably. Finally, Figs. 1 and 4 compare ground truth poses, raw KISS-ICP poses, and corrected trajectories for MSA and the entropy-based method post optimization. Across all illustrated trajectories, the proposed approach achieves consistently lower translation and rotation RPE than other methods, with loop closure segments closely aligning with the ground truth, demonstrating its ability to retain essential edges.

### v3

@@ III-E3 Dynamic Keyframe Selection
To minimize the keyframe set size in real-time, we must consider the entire set $\mathbf{K}$; however, this is impractical for maintaining computational feasibility in loop closure detection. The challenge is to determine the contribution of each keyframe, which requires anticipating future keyframes and introduces a lack of causality. Moreover, finding the optimal subset $\mathbf{K}_{\text{S}}^{*}$ is an NP-hard problem, involving an exhaustive search of all keyframe combinations, which becomes computationally infeasible for sets larger than 15–20 keyframes. The goal of this research is to show that reducing the number of keyframes through an informed sampling strategy can significantly decrease the search space for loop closure detection, improving computational efficiency by reducing memory usage and query time without compromising SLAM accuracy. The proposed approach seeks to balance computational complexity and optimization precision, ensuring scalability and robustness.
IV Minimal Subset Approach
@@
To address these challenges, we propose the Minimal Subset Approach (MSA) to approximate a solution to the problems in Eq. (5)-(6). The proposed MSA identifies and eliminates redundancy within a keyframe set while preserving essential data. Designed for real-time use, it employs a sliding window combinatorial optimization with dual objectives: balancing information preservation and redundancy minimization.
IV-A Redundancy in Keyframes
@@
$\bar{\mathcal{L}}\equiv\mathcal{L}=\big\{\langle i,j\rangle\mid f(\mathbf{d}_{i},\mathbf{d}_{j})>\tau,\;\forall\langle\mathbf{k}_{i},\mathbf{k}_{j}\rangle\in\bar{\mathbf{K}}\times\bar{\mathbf{K}}\big\},$
@@
[…]
@@
IV-B Information Preservation in Keyframes
@@
Let $F:\mathbb{R}^{3}\rightarrow\mathbb{R}^{M}$, represent the descriptor extraction function, either learning-based [ma2022overlaptrans, cattaneo_lcdnet_2022] or handcrafted [Giseop2022scpp, Xu2023ring], which maps each observation from 3-D space to an $M$-dimensional representation. This function does not directly depend on the pose $\mathbf{x}_{t}$, but rather on the input observation $\mathbf{z}_{t}$ (e.g. LiDAR scan), which itself depends on the pose through the sensor model $\mathbf{z}_{t}\propto h(\mathbf{x}_{t},\mathbf{W})$, where $\mathbf{W}$ represents the world. To understand how the descriptors are sensitive to pose changes, we can compute the Jacobian $\mathbf{J}$ of $F$ with respect to poses $\mathbf{x}$ through the chain rule $(\partial\mathbf{F}\,\partial\mathbf{z})/(\partial\mathbf{z}\,\partial\mathbf{x})$. Explicitly deriving the Jacobian can be impractical because of the intricate nature of functions like deep neural networks. Consequently, we employ numerical approximations to estimate the rate of change. Moreover, since poses are encompassed within $SE(3)$, we reduce the dimensional complexity by using the Euclidean norm for distance measurement instead of individual axis derivatives. It is essential to use yaw-invariant descriptors, as suggested in [ma2022overlaptrans, Giseop2022scpp], to ensure the Jacobian calculations are meaningful by preventing orientation changes. Considering the descriptors $\mathbf{d}\in\mathbb{R}^{M}$ as random variables and poses $\mathbf{x}\in SE(3)$ as samples, the product $\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}$ estimates the information matrix: $\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}=({\partial\mathbf{F}}/{\partial\mathbf{x}})^{\mathsf{T}}({\partial\mathbf{F}}/{\partial\mathbf{x}})=\mathbf{V}\mathbf{\Lambda}{\mathbf{V}^{-1}}$, where $\mathbf{\Lambda}$ is the diagonal matrix of eigenvalues and $\mathbf{V}$ is the matrix of eigenvectors from the decomposition. The eigenvectors represent the principal directions of variation, and the eigenvalues represent their magnitudes. The descriptors are transformed using the eigenvectors and scaled by the square root of the eigenvalues, $\mathbf{D}^{\prime}=\sqrt{\mathbf{\Lambda}}\cdot\mathbf{V}\cdot\mathbf{D}$, aligning them with the main directions of variability. Therefore, we can define the information preservation term for a keyframe set as:
$\pi_{\tau}(\mathbf{K})=\frac{1}{N}\sum_{i=0}^{N-1}f_{\delta}(\mathbf{d}_{i}^{\prime},\mathbf{d}_{i+1}^{\prime}),$ (10)
@@
where $0<\pi_{\tau}(\mathbf{K})\leqslant 1$, $f_{\delta}$ is the distance function between descriptors, and $\mathbf{d}^{\prime}\in\mathbf{D}^{\prime}$ are the transformed descriptors. Higher values of $\pi_{\tau}$ indicate that the poses better preserve the variability in the descriptor space. Eigenvectors $\mathbf{v}_{k}$ denote directions of maximal variability, and the $k$-th eigenvalue $\lambda_{k}$ quantifies the variance described by each eigenvector. Larger eigenvalues indicate more significant patterns of variability.
@@
Figure 2: Redundant keyframes effect. Example of removing redundant keyframes (light blue). The updated odometry edge set, $\mathcal{E}_{o}^{*}$, influences the pose graph optimization, while the sparsified loop closure edges, $\bar{\mathcal{L}}=\mathcal{L}\>\backslash\langle 5,7\rangle$, can further reduce the computational complexity.
IV-C Sliding Window Optimization
To manage the computational complexity of optimizing a large keyframe set and proactively selecting keyframes for future queries, we introduce a sliding window optimization method. Figure 3 illustrates the interaction between the front-end and back-end systems, highlighting the contribution of the proposed keyframe sampling strategy in linking the two. This approach continuously optimizes keyframe sets $\mathbf{K}_{t}$ over the mission time $T$, approximating the optimal keyframe set as:
@@ IV-C Sliding Window Optimization
The process begins by initializing a window keyframe set $\mathbf{K}_{t}$ containing $N$ keyframes. The time step $t$ advances once $N$ new keyframes are available, and the optimization gets the optimal window keyframe set $\mathbf{K}_{t}^{*}$. The last selected keyframe becomes the first in the next window, and any previously unselected keyframes are re-evaluated in the context of the newly incoming keyframes. To find this optimum, we generate all possible subsets of $\mathbf{K}_{t}$, forming the power set $\mathbb{P}(\mathbf{K}_{t})$ with cardinality $|\mathbb{P}(\mathbf{K}_{t})|=2^{N}$. To reduce computational complexity, we apply constraints on the power set, retaining only subsets that satisfy min and max distances between consecutive poses, as per Eq. (8), reducing the power set by 5–10 times. The constrained power set can be denoted as:
$\bar{\mathbb{P}}(\mathbf{K}_{t})=\big\{\mathbf{K}_{t}^{\mathbb{S}}\in\mathbb{P}(\mathbf{K}_{t})|\delta_{l}\leqslant\|\mathbf{x}_{i}-\mathbf{x}_{i+1}\|_{2}\leqslant\delta_{u},\forall\mathbf{k}\in\mathbf{K}_{t}^{\mathbb{S}}\big\},$ (12)
where $\delta_{l}$ and $\delta_{u}$ are the lower and upper bounds on inter-pose distance ($\delta_{l}$ prevents near-duplicate/stationary frames; $\delta_{u}$ keeps pairs within descriptor locality). We adaptively select these limits based on the average frame-to-frame distance $\delta_{\text{avg}}$ within a window as $\delta_{l}=\lambda_{l}\cdot\delta_{\text{avg}}$ and $\delta_{u}=\lambda_{u}\cdot\delta_{\text{avg}}$, where $\lambda_{l}$ and $\lambda_{u}$ are constant hyperparameters, further discussed in Section V-C. The window optimization can be formulated as:
$\mathbf{K}_{t}^{*}=$ $\operatorname*{arg\,min}_{\mathbf{K}_{t}^{\mathbb{S}}}\>\>\big({\rho_{\tau}\,(\mathbf{K}_{t}^{\mathbb{S}})+\alpha}\big)\big/\big({\pi_{\tau}(\mathbf{K}_{t}^{\mathbb{S}})}+\beta\big)$ (13)
@@
The minimization problem is solved through an exhaustive search, evaluating each subset’s information matrix $\mathbf{J}_{\mathbf{F}}^{\mathsf{T}}\mathbf{J}_{\mathbf{F}}$ and quantifying the information preservation and redundancy terms, selecting the subset with the best combined score.
Figure 3: Overall pipeline. Block diagram of the overall pipeline, highlighting the interplay between front-end (descriptor extraction and odometry modules) and back-end optimization (place recognition, loop closure detection and pose graph optimization), facilitated by the proposed keyframe sampling scheme.
V Experimental Setup and Results
TABLE I: Translational (t) and Rotational (R) Absolute Trajectory Error Improvement after PGO and the Area Under the PR Curve
@@ V Experimental Setup and Results
Metrics t. ATE [%] R. ATE [%] PR-AUC [%] t. ATE [%] R. ATE [%] PR-AUC [%] t. ATE [%] R. ATE [%] PR-AUC [%]
ScanContext
All Samples 72.7 / 76.1 94.4 / 49.0 99.8 / 99.4 30.2 / 69.4 24.1 / 56.3 99.6 / 99.1 92.0 / 93.7 93.7 / 92.3 74.6 / 91.0
Const. 1.0m 90.5 / 76.2 95.9 / 52.6 97.1 / 96.6 63.2 / 88.6 55.2 / 70.7 96.1 / 85.9 84.3 / 92.5 87.4 / 92.0 62.5 / 89.8
Const. 2.0m 86.2 / 62.3 94.0 / 46.4 71.8 / 78.2 72.1 / 83.3 69.4 / 72.0 79.8 / 83.4 66.7 / 79.2 59.5 / 78.6 51.2 / 81.3
Const. 3.0m 79.8 / 60.1 77.4 / 42.3 74.2 / 74.5 58.7 / 79.5 60.6 / 69.2 64.2 / 73.5 54.4 / 77.1 40.9 / 69.8 41.0 / 68.2
Entropy-based 66.9 / 46.7 93.4 / 48.3 92.0 / 90.9 68.7 / 71.8 69.0 / 66.8 89.1 / 82.8 90.1 / 80.3 89.8 / 82.9 72.5 / 85.8
Spaciousness 77.8 / 55.6 87.6 / 46.7 76.5 / 77.8 67.9 / 70.7 68.5 / 66.1 81.0 / 86.3 79.7 / 79.6 84.2 / 82.1 45.5 / 79.3
MSA (Ours) 90.3 / 76.4 96.0 / 53.2 99.7 / 94.0 71.9 / 91.8 70.0 / 76.8 95.3 / 85.7 91.6 / 91.5 91.4 / 90.2 74.3 / 85.6
OverlapTrans
All Samples 75.7 / 78.6 95.9 / 51.1 99.9 / 99.5 29.8 / 67.4 24.9 / 59.5 99.9 / 89.8 96.2 / 98.1 95.3 / 97.5 75.5 / 90.2
Const. 1.0m 94.6 / 78.7 97.3 / 55.3 97.3 / 96.7 59.3 / 87.1 56.4 / 71.0 96.6 / 86.7 85.8 / 97.1 84.1 / 97.2 62.8 / 89.9
Const. 2.0m 84.4 / 44.9 97.2 / 23.7 72.1 / 78.0 70.0 / 80.9 68.4 / 73.7 80.7 / 83.9 06.8 / 75.9 06.8 / 77.2 50.7 / 81.4
Const. 3.0m 90.5 / 22.6 87.9 / 13.8 74.3 / 74.2 63.6 / 87.7 64.6 / 79.1 64.1 / 72.9 00.0 / 83.6 00.0 / 84.4 41.3 / 67.9
Entropy-based 67.9 / 24.1 93.0 / 09.1 92.2 / 89.7 69.2 / 68.7 69.5 / 63.1 89.4 / 83.8 77.6 / 69.4 76.2 / 72.6 72.1 / 86.0
Spaciousness 85.8 / 35.5 97.2 / 39.1 76.4 / 77.5 68.4 / 68.6 68.1 / 63.1 81.4 / 86.0 03.5 / 67.9 03.5 / 71.7 42.7 / 79.9
MSA (Ours) 93.1 / 78.8 97.6 / 55.5 99.8 / 93.8 70.2 / 91.2 69.9 / 77.5 95.5 / 86.3 93.0 / 98.1 92.0 / 97.5 75.0 / 85.2
@@
The bold results correspond to the best performing sampling method, while the underlined results correspond to the second best performing sampling method.
@@
To support the claims of our contributions, on how the sampling interval affects both place recognition and pose graph optimization performance, we conducted a series of experiments with the following setup. Descriptor extraction was performed using two approaches: the learning-based OverlapTransformer (OT[ma2022overlaptrans], which generates a $1\times 256$ feature vector, and the hand-crafted Scan Context (SC[Giseop2022scpp], which produces a $20\times 60$ feature representation. These choices demonstrate the descriptor-agnostic nature of our method and its compatibility with both learned and hand-crafted descriptors. We benchmark our approach against constant sampling intervals of 1, 2, and 3 meters, commonly found in algorithms such as LIO-SAM [liosam2020shan], LAMP [Chang2022lamp2] and DiSCo-SLAM [Huang2022disco], as well as adaptive sampling strategies based on LiDAR spaciousness [chen2022dlo] and entropy [Zeng2023entropy]. All experiments were conducted on a 14th Gen Intel Core i9 system with 128GB of RAM, where the sampling window optimization ($N=10$) is solved in approximately 5–15 milliseconds, making it suitable for real-time sampling.
@@
V-A Quantitative and Qualitative Analysis
For the batch experiments in Table I, we evaluate our approach on three datasets: KITTI Odometry [Geiger2012CVPR], MulRan [Gisep2020mulran], and Apollo-SouthBay [Weixin2019apollo], representing diverse environments, including urban areas, rural landscapes, and complex structures like bridges and tunnels. Keyframes are sampled for all methods, and the pose graph is constructed using GTSAM [gtsam] with odometry provided from KISS-ICP [vizzo2023ral]. Loop closure candidates are identified using OT or SC descriptors with a similarity threshold of 0.8. Candidates within a 3-meter radius of the ground truth are classified as true positives and processed with small_gicp [small_gicp] to estimate relative transformations. Matches with registration residuals below 0.3 meters are verified and added as edges to the pose graph, which is optimized using the Levenberg-Marquardt algorithm.
Table I summarizes the results for all datasets, descriptors, and methods. The metrics include the Absolute Trajectory Error (ATE) for both translation and rotation, presented as the percent improvement in the KISS-ICP trajectory after pose graph optimization with each method. Additionally, we report the Area Under the Precision-Recall Curve (PR-AUC), as we demonstrate that sampling impacts place recognition performance. The results demonstrate that the proposed MSA consistently delivers the best performance, while constant intervals yield varying results across different environments, highlighting the absence of a universal fixed sampling interval. In Fig. 4, the data from Table I are presented as box plots, illustrating the deviation of each method’s ATE performance compared to using all samples, alongside variations in memory usage and total pipeline execution time. The proposed approach sustains overall improvements with minimal performance loss in both translation and rotation, while significantly reducing memory usage and execution time. Although other methods also reduce memory and processing overhead, their performance lacks robustness and varies considerably. Finally, Figs. 1 and 5 compare ground truth poses, raw KISS-ICP poses, and corrected trajectories for MSA and the entropy-based method post optimization. Across all illustrated trajectories, the proposed approach achieves consistently lower translation and rotation RPE than other methods, with loop closure segments closely aligning with the ground truth, demonstrating its ability to retain essential edges.